OpenStudy (anonymous):

How do you write 0.3 as a fraction? 1/4 3/5 4/9 1/3

OpenStudy (anonymous):

0.3 = 3/10

OpenStudy (zarkon):

based on their last question, I believe they mean \(.\overline{3}\)

OpenStudy (anonymous):

probably

OpenStudy (anonymous):

so in that case what is the answer?

OpenStudy (anonymous):

.3=3/10 = 30% (about) = 1/3

OpenStudy (anonymous):

thanks